OpenStudy (anonymous):

If 3 times a number is added to -9, the result equals to 12 times the number, find the number?

OpenStudy (a_clan):

3x-9=12x x=-1

OpenStudy (anonymous):

3n-9=12n -9=9n n=-1

jhonyy9 (jhonyy9):

-9+3x=12 3x=21 x=21/3 =7 x=7